# Taming Data and Transformers for Audio Generation

### v1

@@ 3.1 Automatic Audio Captioning
To avoid degrading the quality of the pretrained BART and audio encoder models, we adopt a two-stage training procedure. In Stage 1, both the audio encoders and BART model are kept frozen, thus allowing the Q-Former, projection layers, and newly introduced delimiter tokens to align to the existing BART input representation. In this stage, we pretrain the model using a larger dataset of weakly-labeled audio clips. In Stage 2, we unfreeze all BART model parameters apart from the embedding layer and finetune the model on Audiocaps dataset at a lower learning rate to make the captioning style align more closely to the target dataset. This training strategy effectively leverages the larger, weakly-labeled dataset while minimizing the knowledge drift in the pretrained BART.
3.2 Data Selection and Re-captioning
Generative models in the image and video domains have shown benefits from increased quantities of data and improved quality of captions. In the audio domain, however, the major human-annotated audio-text datasets, namely AudioCaps [46] and Clotho [18], comprise a total of only 51k audio clips combined. Leveraging our trained AutoCap model, we aim to significantly expand the number of such videos by pairing them with high-quality synthetic captions.
We start dataset construction by selecting the AudioSet [28] and VGG-Sound [4] datasets as ideal candidates for re-captioning as each audio clip is derived from a video with an associated title and description. Additionally, since these datasets were constructed to be audio-oriented, they tend to have good alignment between the visual and audio components. To enrich the metadata information, we consider captions derived from the video modality using the video captioning method of Chen et al[8]). In addition, we filter AudioSet to exclude content labeled as ‘music’ and ‘speech’ as we are more interested in the generation of audio effects. We sample 10-second audio clips randomly from each selected video. Furthermore, we select Freesound, BBC Sound Effects, and SoundBible as additional data sources. We input their associated title to the audio captioning model and leave the additional video caption metadata empty as no video modality is present. Tab. 1 summarizes the data sources employed to build our dataset. In total we collect and re-caption 761,000 audio clips, forming the largest available dataset of audio with paired captions.
We report in Tab. 2 a comparison of our collected dataset with the most popular audio text datasets. Our dataset possesses the largest number of audio clips and the longest total duration. Unlike WavCaps, whose captions are produced using the raw data and without considering the audio modality, our dataset is captioned by leveraging additional metadata derived from the video modality while maintaining an audio-centric approach. Compared to LAION-Audio-630k, the largest available dataset in terms of text-audio pairs, we offer an increased number of audio clips and more than double the total duration. Moreover, our dataset presents higher quality captions compared to LAION-Audio captions which are directly acquired from the raw data and thus present a high level of noise.
Data Source Num. of pairs
AudioSet 339,387
VGGSounds 126,905
Freesounds 262,300
BBC Sound Effects 31,201
SoundBible 1,320
Our dataset 761,113
Table 2: Comparative overview of the main audio-language datasets.
Dataset # Text-Audio Pairs Duration (h) Text source
AudioCaps 52,904 144.94 Human
Clotho 5,929 37.00 Human
MACS 3,537 9.83 Human
WavText5K 4,072 23.20 Online raw-data
SoundDescs 32,979 1,060.4 Online raw-data
LAION-Audio-630K 633,526 4,325.39 Online raw-data
WavCaps 403,050 7,567.92 Processed raw-data
Ours 761,113 8,763.12 Automatic re-captioning
@@ 3.3 Scalable Text-2-Audio Generation
We design our audio generation pipeline, GenAu, as a latent diffusion model. Fig. 1 (right) shows an overview of our proposed model. In the following section, we describe in detail the structure of our latent variational autoencoder (VAE) and the latent diffusion model.
@@
Latent VAE. Directly modeling waveform audio data is complex due to the high data dimensionality of audio signals. Instead, we replace the waveform with a Mel-spectrogram representation and use a VAE to further reduce its dimensionality, following prior work [79, 43]. We note that the once generated, a Mel-spectrogram representation can be decoded back to a waveform through the use of an audio vocoder [50]. However, commonly-used 2D autoencoder designs [65, 66, 79], are not well suited to the Mel-spectrogram representation, as the separation between the Mel channels is non-linear and thus not well suited for 2D convolutions. We instead opt for a 1D-VAE design based on 1D convolutions similar to a recent work [42], with the additional benefit of a more compact audio representation. We train our 1D-VAE model using a combination of reconstruction, adversarial, and KL regularization losses following [22].
@@
Latent diffusion model. Following the latent diffusion paradigm, we generate audio by training a diffusion model in the latent space of the 1D-VAE. Transformer models currently attain state-of-the-art performance in audio generation [42]. To improve model scalability, we propose to use an efficient transformer architecture due to its success in handling long-range interactions as in video generation [7, 80]. In particular, we adopt the FIT architecture of Menapace et al[80] which was originally proposed to work in the pixel space, and revise it for the latent space of the audio modality.
@@
Given a 1D input $\mathbf{x}$, we first apply a projection operation to produce a sequence of input patch tokens. We then apply a sequence of FIT blocks to the input patches where each block divides patch tokens into contiguous groups of a predefined size. A set of local self-attention layers are then applied separately to each group to avoid the quadratic computational complexity of attention computation. Differently from the video domain [80] where the high input dimensionality makes the local layers excessively expensive, we found them to be beneficial for audio generation. To further reduce the amount of computation while maintain long-range interaction, each block considers a small set of latent tokens. First, a read operation implemented as a cross-attention layer transfers information from the patches to the latent tokens. Later, a series of global self-attention operations are applied to the latent tokens, allowing information-sharing between different groups. Finally, a write operation implemented as a cross-attention layer transfers information from the latent tokens back to the patches. Due to the reduced number of latent tokens when performing the global self-attention, computational requirements of the model are reduced with respect to a vanilla transformer design [109].
@@
To condition the generation on an input prompt, we use a pretrained FLAN-T5 model [13] and a CLAP [20] text encoder to produce the their respective embeddings $e_{\mathrm{FLAN}}$ and $e_{\mathrm{CLAP}}$, which we concatenate with the diffusion timestep $t$ to form the input conditioning signal $c$. We insert an additional cross attention operation inside each FIT block immediately before the ‘read’ operation that makes latent tokens attend to the conditioning. Moreover, we use a conditioning on dataset ID to adapt the generation style to different types of datasets.
@@
We follow a linear noise scheduler and train the model using the epsilon diffusion objective:
$\mathcal{L}=\mathbb{E}_{t,\mathbf{x},\boldsymbol{\epsilon}}~{}\big{\lVert}\mathcal{G}(\mathbf{x}_{t},c)-\boldsymbol{\epsilon}\big{\rVert}^{2}_{2}\text{,}$ (6)
@@
where $\mathcal{G}$ is the FIT generator backbone, $\mathbf{x}_{t}$ is the input with applied noise at diffusion timestep $t$, and $\boldsymbol{\epsilon}$ is noise sampled in $N(0,1)$ with the same shape as the input.
4 Experiments
We structure the experiments section as follows: Sec. 4.1 evaluates AutoCap by quantitatively comparing it to previous work. Sec. 4.2 demonstrates the capabilities of GenAu quantitatively. For both, we discuss training details, baselines, metrics, results, and ablations.
4.1 Automatic Audio Captioning
Training dataset and details. We train our captioning model in two stages. During stage 1, we pretrain on a large weakly labeled dataset of 634,208 audio clips, constructed from AudioSet [28], Freesound, BBC Sound Effects, SoundBible, AudioCaps [46], and Clotho [18] datasets. We use the ground truth captions from AudioCaps and Clotho dataset, WavCaps captions for Freesound, SoundBible, and BBC Sound Effects, and handcrafted captions through a template leveraging the provided ground truth class labels for AudioSet. We use the title provided with each clip, and pre-compute video captions using a pretrained Panda70M model [8] for the clips that are associated with video modality and pass an empty string otherwise. We pretrain the model for 20 epochs with a learning rate of 1e-4. We keep the audio encoder and pretrained BART frozen during this stage. In Stage 2, we fine-tune the model for 20 epochs only on AudioCaps using a learning rate of 1e-5. We randomly sample 10-second clips at 32KHz for all of our captioning experiments.
Baselines. We compare with ACT [76], V-ACT [71], BART-tags [31], AL-MixGEN [47], ENCLAP [49], HTSAT-BART [123] and CNext-trans [136]. Among these baselines, ENCLAP and CNext-trans achieve the best performance. ENCLAP benefits from a stronger audio encoder and the use of a CLAP representation for additional guidance. CNext-trans trains a lightweight transformer instead of fine-tuning a pretrained language model to reduce overfitting.
Metrics and evaluation. We report results using the the established BLEU1 [86], BLEU2 [86], ROUGE [63], Meteor [56], CIDEr [110], and SPIDEr [69] metrics. We evaluate our method on the AudioCaps test split using the last checkpoint. We use only 876 clips as some videos were deleted since the original data release. We follow the same evaluation pipeline as baselines and include their reported results. Results that were not provided in these publications are excluded from our analysis.
@@ 4.1 Automatic Audio Captioning
ACT [76] AS 64.7 25.2 46.8 22.2 67.9 16.0 42.0
V-ACT [71] - 69.8 28.1 49.4 23.7 71.1 17.2 44.2
BART-tags [31] AS 69.9 26.6 49.3 24.1 75.3 17.6 46.5
AL-MixGEN [47] - 70.0 28.9 50.2 24.2 76.9 18.1 47.5
ENCLAP-Large [49] - - - - 25.5 80.2 18.8 49.5
HTSAT-BART [123] - 67.5 27.2 48.3 23.7 72.1 16.9 44.5
HTSAT-BART [123] AC+CL+WC 70.7 28.3 50.7 25.0 78.7 18.2 48.5
CNext-trans [136] - - - - - - - 46.6
CNext-trans [136] AC+CL+MA+WC - - - 25.2 80.6 18.4 49.5
@@
Table 4: Ablation study on AudioCaps through evaluating key model components.
@@
- Unfreeze Word Embedding 22.5 82.6 18.1 50.4
Results. In Tab. 3 we report the quantitative comparison. Our method outperforms previous methods on all metrics, achieving notable improvements in the CIDEr and BLUE1 scores, with values of 83.2 and 73.1, respectively. We found that incorporating metadata significantly enhances the CIDEr scores but slightly reduces the SPICE scores. This trade-off likely results from the enhanced descriptive detail brought by the metadata, which, while enriching the content, introduces noise that may compromise the model’s semantic precision. In addition, AudioCaps is labeled based on audio information alone, thus evaluation penalizes the description of information that can not be deduced with certainty from the audio modality only, such as the specific type of object producing a rustling sound. Compared to ENCLAP-Large [49], and CoNeTTE (CNext-trans) [136], we find the captions produced by our method to be more descriptive and precise with a better temporal understanding. ENCLAP-Large often misses important details and exhibits lower temporal accuracy. CNext-trans, while accurate, often produces short captions that lack details. We include qualitative comparisons in the project Website. Moreover, AutoCap is four times faster than ENCALP, producing a caption for a 10-second clip in 0.28 seconds, compared to ENCALP which takes 1.12 seconds. Furthermore, we observe consistent improvements when pretraining on a large scale of weakly-labeled data during first stage, validating the effectiveness of our training strategy in benefiting from a larger, weakly-labeled dataset.
Ablations. In Tab. 4, we ablate model design choices. We observe the use of the CLAP embedding to bring a 2.5 points increase in the CIDEr score. We also validate that when not performing Stage 2 training, which involves finetuning of the BART [58] model, performance degrades on all metrics, a finding we attribute to the necessity of adapting BART’s decoder to the sentence structure typical of AudioCaps. A more severe degradation in performance is observed if Stage 1 is not performed, with the misaligned representation between the encoder and the decoder causing catastrophic forgetting in the language model. Finally, if BART word embeddings are finetuned in Stage 2 instead of being kept frozen, we observe a slight performance degradation.
@@ 4.2 Text-2-Audio Generation
Model Prams FD $\downarrow$ IS $\uparrow$ FAD $\downarrow$ $\text{CLAP}_{LAION}$uparrow $\text{CLAP}_{MS}\uparrow$
GroundTruth - - - - 0.251 0.671
AudioLDM-L [65] 739M 37.89 7.14 5.86 - 0.429
AudioLDM 2-L [66] 712M 32.50 8.54 5.11 0.212 0.621
TANGO [29] 866M 26.13 8.23 1.87 0.185 0.597
TANGO 2 [74] 866M 19.77 8.45 2.74 0.264 0.590
Make-An-Audio [43] 453M 27.93 7.44 2.59 0.207 0.621
Make-An-Audio 2 [42] 937M 15.34 9.58 1.27 0.251 0.645
GenAu-Small 493M 15.54 11.10 1.07 0.271 0.674
GenAu-Large 1.25B 16.51 11.75 1.21 0.285 0.668
Table 6: GenAu ablation results on AudioCaps. Training is performed on AudioCaps only, or on our full dataset with or without AutoCap re-captioning.
Model Training Data Re-Captioning FD $\downarrow$ FAD $\downarrow$ IS $\uparrow$
GenAu-Small AC N/A 15.96 1.21 9.55
GenAu-Large AC N/A 16.89 1.51 10.25
GenAu-Small Full 17.02 1.85 10.52
GenAu-U-Net Full 25.57 1.98 9.54
GenAu-Small Full 15.54 1.07 11.10
GenAu-Large Full 16.51 1.21 11.75
Training dataset and details. We use our best-performing captioning model to re-caption the WavCaps dataset. In addition, we obtain 339,387 videos from AudioSet and 126,905 videos from VGGSounds, totaling 761,113 clips. For those obtained from sound-only platforms, we input an empty string as the video caption. For full details of the data sources of our training dataset, please refer to Tab. 1. We additionally use Clotho and AudioCaps training datasets with their ground truth caption. To stay consistent with baselines, we train at 16kHz resolution. We use a patch size of 1 and a group size of 32. We use LAMB optimizer [129] with a LR of 5e-3. We train for 220k steps and choose the checkpoint with the highest IS, at step 210k and 207k for the large and small model. We also disable EMA as found it to make the metrics unstable.
Baselines. We compare with TANGO 1 & 2, [29], AudioLDM 1 & 2 [65, 66], and Make-An-Audio 1 & 2 [43, 42]. Both AudioLDM and Make-an-Audio train a UNet-based latent diffusion models [92] on Mel-Spectrogram representation of the audio, by regarding the Mel-Spectrogram as a single channel image, and use a pretrained CLAP encoder to condition the generation on an input prompt. TANGO proposed to use FLAN-T5 [13] as the text encoder and reported significant improvements. AudioLDM-2 and Make-an-Audio-2 proposed to use a dual encoder strategy of a T5 [90] and CLAP encoder. AudioLDM-2 focused on extending the generation and conditioning to various domains. Specifically, they use language of audio (LOA) to condition the generation on images, audio, or transcripts and train their model for music and speech generation. Make-an-Audio-2 proposes to use a 1D VAE representation and employ a feed-forward Transformer-based model to replace the UNet. Recently, Tango-2 proposed to use instruction fine-tuning on a synthetic dataset to enhance the temporal understanding. In our experiments, we focus on text-conditioned natural audio generation and generate 10s clips at a resolution of 16Khz. We only provide a qualitative comparison with StableAudio [24] in the webpage since it reports different metrics without any checkpoints.
@@
Metrics. We compare the performance of our method with baselines using the standard Frechet Distance (FD), Inception score (IS), and CLAP score on Audioset test dataset, containing 964 samples. These metrics use pretrain PANNs [51] model features. Some prior work also reported the Frechet Distance results using the VGGish network [39], denoted as (FAD) [45]. Additionally, to compute the CLAP score, some prior work [66] used CLAP from LAION [120], while others [74, 43, 42] used CLAP from Microsoft [20]. To avoid confusion, we report both metrics and denote them as $\text{CLAP}_{LAION}$ and $\text{CLAP}_{MS}$, respectively. We follow the same evaluation protocols of AudioLDM [65] and use the AudioLDM evaluation package to compute the metrics. Due to inconsistencies in evaluation pipelines and varying results for the same baselines reported in different studies, we have chosen to recompute all metrics using the official checkpoints to ensure consistent comparisons.

### v2

@@ 3.1 Automatic Audio Captioning
To avoid degrading the quality of the pretrained BART and audio encoder models, we adopt a two-stage training procedure. In Stage 1, both the audio encoders and BART model are kept frozen, thus allowing the Q-Former, projection layers, and newly introduced delimiter tokens to align to the existing BART input representation. In this stage, we pretrain the model using a larger dataset of weakly-labeled audio clips. In Stage 2, we unfreeze all BART model parameters apart from the embedding layer and finetune the model on the Audiocaps dataset at a lower learning rate to make the captioning style align more closely to the target dataset. This training strategy effectively leverages the larger, weakly-labeled dataset while minimizing the knowledge drift in the pretrained BART. The use of Qformer to learn an intermediate representation is pivotal for such training strategy. Furthermore, the Qformer summarizes the audio representation into four times fewer tokens, significantly reducing the inference time.
Figure 2: Audio data collection pipeline. We employ online video transcripts to identify audio segments without speech or music. These are processed by AutoCap to generate captions. We retain only ambient clips with captions lacking music and speech keywords.
3.2 Data Collection and Re-captioning Pipeline
Generative models in the image and video domains have shown benefits from increased quantities of data and improved quality of captions. In the audio domain, however, the major human-annotated audio-text datasets, namely AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020), provide only 51k audio clips combined. Previous methods attempted to extract additional ambient audio clips from existing video datasets using pretrained audio classifiers, but a high rejection rate of $99\%$ marked this method impractical. Instead, we found that automatic transcripts offer reliable information about the segments containing ambient sounds. In particular, we propose to select the parts of the videos that contain no automatic transcription, suggesting the absence of speech and music. Such an approach offers specific advantages over using pretrained classifiers. Automatic transcripts, readily available for most online videos, eliminate the need to download and process video and audio data before filtering. Additionally, as these transcripts provide precise time-aligned information, they facilitate the extraction of more segments, effectively reducing the rejection rate to $83\%$. Subsequently, we leverage our AutoCap model to provide textual descriptions of the extracted audio clips. Despite the effectiveness of this method in collecting ambient sounds, some clips still inadvertently contain music or speech due to transcription errors, particularly with speech in less common languages. We address this by analyzing captions and filtering out clips with keywords related to speech or music. Figure 2 summarizes our data collection pipeline.
We follow this process to extract 466k audio-text pairs from Audioset (Gemmeke et al., 2017) and VGGSounds (Chen et al., 2020). Additionally, we recaption audio-only dataset such as Freesound, BBC Sound Effects, and SoundBible. To provide metadata, we employ the captioning model of Chen et al. (2024b) to extract a caption whenever a video content is available and pass an empty text otherwise. In total, we form AutoReCap, a large-scale dataset compromising of 761,113 audio-text pairs with precise captions. As an additional contribution, we introduce AutoReCap-XL, in which we scale our approach by analysing four additional large-scale video dataset (Lee et al., 2021; Xue et al., 2022; Zellers et al., 2022; Nagrani et al., 2022) with a total of 71M videos and $715.4k$ hours. In total, we collect and re-caption 57M ambient audio clips spanning $123.5k$ hours from 20.3M different videos, forming by far the largest available dataset of audio with paired captions. More details about the dataset can be found in the Appendix.
Figure 3: Scaling analysis of model size (left) and data volume with synthetic captions (right) reveal consistent improvements in FD and IS.
@@ 3.3 Scalable Text-2-Audio Generation
We design our audio generation pipeline, GenAu, as a latent diffusion model. Figure 1 (right) shows an overview of our proposed model. In the following section, we describe in detail the structure of our latent variational autoencoder (VAE) and the latent diffusion model.
@@
Latent VAE. Directly modeling waveform audio data is complex due to the high data dimensionality of audio signals. Instead, we replace the waveform with a Mel-spectrogram representation and use a VAE to further reduce its dimensionality, following prior work (Melechovsky et al., 2024; Huang et al., 2023b). Once generated, Mel-spectrograms can be decoded back to a waveform through the use of an audio vocoder (Kong et al., 2020). However, commonly-used 2D autoencoder designs (Liu et al., 2023b; c; Melechovsky et al., 2024), are not well suited to the Mel-spectrogram representation, as the separation between the Mel channels is non-linear, which is not well suited for 2D convolutions. We instead opt for a 1D-VAE design based on 1D convolutions similar to Huang et al. (2023a). We train our VAE using a combination of reconstruction, adversarial, and KL regularization losses following Esser et al. (2021).
@@
Latent diffusion model. Following the latent diffusion paradigm, we generate audio by training a diffusion model in the latent space of the 1D-VAE. Transformer-based diffusion models currently attain state-of-the-art performance in audio generation (Huang et al., 2023a). To improve model scalability, we propose to use an efficient transformer architecture due to its success in handling long-range interactions as in video generation (Chen & Li, 2023; Menapace et al., 2024). In particular, we adopt the FIT architecture of Menapace et al. (2024) which was originally proposed to work in the pixel space and revise it for the latent space of the audio modality.
@@
Given a 1D input $\mathbf{x}$, we first apply a projection operation to produce a sequence of input patch tokens. We then apply a sequence of FIT blocks to the input patches where each block divides patch tokens into contiguous groups of a predefined size. A set of local self-attention layers are then applied separately to each group to avoid the quadratic computational complexity of attention computation. Differently from the video domain (Menapace et al., 2024) where the high input dimensionality makes the local layers excessively expensive, we found them to be beneficial for audio generation. To further reduce the amount of computation while maintaining long-range interaction, each block considers a small set of latent tokens. First, a read operation implemented as a cross-attention layer transfers information from the patches to the latent tokens. Later, a series of global self-attention operations are applied to the latent tokens, allowing information-sharing between different groups. Finally, a write operation implemented as a cross-attention layer transfers information from the latent tokens back to the patches. Due to the reduced number of latent tokens when performing the global self-attention, computational requirements of the model are reduced with respect to a vanilla transformer design (Vaswani et al., 2017). Such a design is also particularly suited for the audio modality, which contains mostly silent or redundant parts. Unlike DiT and UNet-based methods (Ronneberger et al., 2015; Peebles & Xie, 2023) which allocate the computation resources uniformly across input tokens, the FiT architecture selectively focuses on the more informative parts.
@@
To condition the generation on an input prompt, we use a pretrained FLAN-T5 model (Chung et al., 2022) and a CLAP (Elizalde et al., 2023) text encoder to produce the their respective embeddings $e_{\mathrm{FLAN}}$ and $e_{\mathrm{CLAP}}$, which we concatenate with the diffusion timestep $t$ to form the input conditioning signal $c$. We insert an additional cross-attention operation inside each FIT block immediately before the ‘read’ operation that makes latent tokens attend to the conditioning. Moreover, we use conditioning on dataset ID to adapt the generation style to different types of datasets.
@@
We follow a linear noise scheduler and train the model using the epsilon prediction objective:
$\mathcal{L}=\mathbb{E}_{t,\mathbf{x},\bm{\epsilon}}~{}\big{\lVert}\mathcal{G}(\mathbf{x}_{t},c)-\bm{\epsilon}\big{\rVert}^{2}_{2}\text{,}$ (6)
@@
where $\mathcal{G}$ is the FIT generator backbone, $\mathbf{x}_{t}$ is the input with applied noise at diffusion timestep $t$, and $\bm{\epsilon}$ is noise sampled in $N(0,1)$ with the same shape as the input.
@@
ACT AS 64.7 25.2 46.8 22.2 67.9 16.0 42.0
V-ACT - 69.8 28.1 49.4 23.7 71.1 17.2 44.2
BART-tags AS 69.9 26.6 49.3 24.1 75.3 17.6 46.5
AL-MixGEN - 70.0 28.9 50.2 24.2 76.9 18.1 47.5
ENCLAP-Large - - - - 25.5 80.2 18.8 49.5
HTSAT-BART - 67.5 27.2 48.3 23.7 72.1 16.9 44.5
HTSAT-BART AC+CL+WC 70.7 28.3 50.7 25.0 78.7 18.2 48.5
CNext-trans - - - - - - - 46.6
CNext-trans AC+CL+MA+WC - - - 25.2 80.6 18.4 49.5
@@
4 Experiments
We structure the experiments section as follows: section 4.1 evaluates AutoCap by quantitatively comparing it to previous work, section 4.2 demonstrates the capabilities of GenAu quantitatively. For both, we discuss training details, baselines, metrics, results, and ablations.
4.1 Automatic Audio Captioning
Training dataset and details. We train our captioning model in two stages. During stage 1, we pretrain on a large weakly labeled dataset of 634,208 audio clips, constructed from AudioSet, Freesound, BBC Sound Effects, SoundBible, AudioCaps, and Clotho datasets. We use the ground truth captions from AudioCaps and Clotho dataset, WavCaps captions for Freesound, SoundBible, and BBC Sound Effects, and handcrafted captions through a template leveraging the provided ground truth class labels for AudioSet. As metadata, we use the title provided with each clip, and pre-compute video captions using a pretrained Panda70M model (Chen et al., 2024b) when the video modality is available or pass an empty string otherwise. We pretrain the model for 20 epochs with a learning rate of 1e-4, while keeping the audio encoder and pretrained BART frozen. In Stage 2, we fine-tune the model for 20 epochs only on AudioCaps using a learning rate of 1e-5. We randomly sample 10-second clips at 32KHz for all of our captioning experiments.
Baselines. We compare with ACT (Mei et al., 2021), V-ACT (Liu et al., 2023e), BART-tags (Gontier et al., 2021), AL-MixGEN (Kim et al., 2022), ENCLAP (Kim et al., 2024b), HTSAT-BART (Xu et al., 2023) and CNext-trans (Étienne Labbé et al., 2023). Among these baselines, ENCLAP and CNext-trans achieve the best performance. ENCLAP benefits from a stronger audio encoder and the use of a CLAP representation for additional guidance. CNext-trans trains a lightweight transformer instead of fine-tuning a pretrained language model to reduce overfitting.
Metrics and evaluation. We report results using the the established BLEU1 (Papineni et al., 2002), BLEU2 (Papineni et al., 2002), ROUGE (Lin, 2004), Meteor (Lavie & Agarwal, 2007), CIDEr (Vedantam et al., 2015), and SPIDEr (Liu et al., 2017) metrics. We evaluate our method on the AudioCaps test split using the last checkpoint of our trained model. We used only 876 clips for evaluation as some videos were deleted since the original data release. We follow the same evaluation pipeline as baselines and include their reported results. Results that were not provided in these publications are excluded from our analysis.
Results. In Tab. 1 we report the quantitative comparison. Our method outperforms previous methods on all metrics, achieving notable improvements in the CIDEr and BLUE1 scores, with values of 83.2 and 73.1, respectively. We found that incorporating metadata significantly enhances the CIDEr scores but slightly reduces the SPICE scores. This trade-off likely results from the enhanced descriptive detail brought by the metadata, which while enriching the content, introduces noise that may compromise the model’s semantic precision. In addition, AudioCaps is labeled based on audio information alone. Thus, the evaluation penalizes the description of information that can not be deduced with certainty from the audio modality only, such as the specific type of object producing a rustling sound. Compared to ENCLAP-Large (Kim et al., 2024b), and CNext-trans (Étienne Labbé et al., 2023), we find the captions produced by our method to be more descriptive and precise with a better temporal understanding. ENCLAP-Large often misses important details and exhibits lower temporal accuracy. CNext-trans, while accurate, often produces short captions that lack details. We include qualitative comparisons in the project Website. Moreover, AutoCap is four times faster than ENCALP, producing a caption for a 10-second clip in 0.28 seconds, compared to ENCALP which takes 1.12 seconds. Furthermore, we observe consistent improvements when pretraining on a large scale of weakly-labeled data during the first stage, validating the effectiveness of our training strategy in benefiting from a larger, weakly-labeled dataset.
Table 2: AutoCap ablation study on AudioCaps.
@@ 4.1 Automatic Audio Captioning
- Unfreeze Word Embed 22.5 82.6 18.1 50.4
Table 3: GenAu ablation study on out-of-distribution dataset.
Model IS FD $\text{CLAP}_{MS}$
GenAU-L 18.98 20.81 0.38
GenAU-L (AC) 12.14 25.82 0.30
GenAU-S 15.76 21.29 0.36
GenAU-S w/o Recap. 11.83 25.34 0.29
Ablations. In Tab. 2, we ablate model design choices. We observe the use of the CLAP embedding to bring a $2.5$ points increase in the CIDEr score. We also validate that when not performing Stage 2 training, which involves finetuning of the BART (Lewis et al., 2020) model, performance degrades on all metrics, a finding we attribute to the necessity of adapting BART’s decoder to the sentence structure typical of AudioCaps. A more severe degradation in performance is observed if Stage 1 is not performed, with the misaligned representation between the encoder and the decoder causing catastrophic forgetting in the language model. Finally, if BART word embeddings are finetuned in Stage 2 instead of being kept frozen, we observe a slight performance degradation.
@@ 4.2 Text-2-Audio Generation
Training dataset and details. We train on similar data settings to baselines. We use our best-performing captioning model to re-caption the WavCaps dataset. In addition, we obtain 339,387 videos from AudioSet and 126,905 videos from VGGSounds, totaling 761,113 clips. For those obtained from sound-only platforms, we input an empty string as the video caption. For full details of the data sources of our training dataset, please refer to the Appendix. We additionally use Clotho and AudioCaps training datasets with their ground truth caption. To stay consistent with baselines, we train at 16kHz resolution. We use a patch size of 1 and a group size of 32. We use LAMB optimizer (You et al., 2020) with a LR of 5e-3. We train for 220k steps and choose the checkpoint with the highest IS, at steps 210k and 207k for the large and small models. We also disable EMA as found it to make the metrics unstable.
Baselines. We compare with TANGO 1 & 2, (Ghosal et al., 2023), AudioLDM 1 & 2 (Liu et al., 2023b; c), and Make-An-Audio 1 & 2 (Huang et al., 2023b; a). Both AudioLDM and Make-an-Audio train a UNet-based latent diffusion model (Rombach et al., 2022) on Mel-Spectrogram representation of the audio, by regarding the Mel-Spectrogram as a single channel image, and use a pretrained CLAP encoder to condition the generation on an input prompt. TANGO proposed to use FLAN-T5 (Chung et al., 2022) as the text encoder and reported significant improvements. AudioLDM-2 and Make-an-Audio-2 proposed to use a dual encoder strategy of a T5 (Raffel et al., 2022) and CLAP encoder. AudioLDM-2 focused on extending the generation and conditioning to various domains. Specifically, they use the language of audio (LOA) to condition the generation on images, audio, or transcripts and train their model for music and speech generation. Make-an-Audio-2 proposes to use a 1D VAE representation and employ a feed-forward Transformer-based model to replace the UNet. Recently, Tango-2 proposed to use instruction fine-tuning on a synthetic dataset to enhance the temporal understanding. In our experiments, we focus on text-conditioned natural audio generation and generate 10s clips at a resolution of 16Khz.
@@
Model Prams ##\## Samples FD $\downarrow$ IS $\uparrow$ FAD $\downarrow$ $\text{CLAP}_{LAION}$uparrow $\text{CLAP}_{MS}\uparrow$
GroundTruth - - - - - 0.251 0.671
AudioLDM-L 739M 634k 37.89 7.14 5.86 - 0.429
AudioLDM 2-L 712M 760k 32.50 8.54 5.11 0.212 0.621
TANGO 866M 45k 26.13 8.23 1.87 0.185 0.597
TANGO 2 866M 60k 19.77 8.45 2.74 0.264 0.590
Make-An-Audio 453M 1M 27.93 7.44 2.59 0.207 0.621
Make-An-Audio 2 937M 1M 15.34 9.58 1.27 0.251 0.645
GenAu w/ U-Net 462M 811K 25.57 9.54 1.98 - -
GenAu-Large 1.25B 811K 16.51 11.75 1.21 0.285 0.668
Table 5: User study between various baselines. % of votes in favor of the baseline to the left.
Model Realism Quality Prompt Alignment Overall Preference
GenAU-L vs GenAU-S 61.20% 58.00% 61.20% 60.40%
GenAU-L vs GenAU-L (AC) 60.40% 54.80% 60.40% 59.20%
GenAU-L vs MAD-2 64.00% 62.40% 68.40% 66.40%
GenAU-S w/o Recap. vs MAD-2 64.40% 64.-0% 63.20% 64.80%
@@
Metrics. We compare the performance of our method with baselines using the standard Frechet Distance (FD), Inception score (IS), and CLAP score on the Audioset test dataset, containing 964 samples. There is little consistency between baselines when computing the metrics. Some prior work reported the Fréchet distance results using the VGGish network (Hershey et al., 2017), denoted as (FAD) (Kilgour et al., 2019), while other uses PANNs (Kong et al., 2019). Additionally, to compute the CLAP score, some prior work (Liu et al., 2023c) used CLAP from LAION, which we denote as $\text{CLAP}_{LAION}$ (Wu et al., 2023), while others (Majumder et al., 2024; Huang et al., 2023b; a) used CLAP from Microsoft (Elizalde et al., 2023), which we denote as $\text{CLAP}_{MS}$. Furthermore, some prior (Liu et al., 2023b; c) used CLAP re-ranking with 3 samples for computing the metrics. Due to such inconsistencies in evaluation pipelines and varying results for the same baselines reported in different studies, we recompute all metrics using the official checkpoints to ensure consistent comparisons. We follow the same evaluation protocols of AudioLDM (Liu et al., 2023b) without CLAP re-ranking and use the AudioLDM evaluation package to compute the metrics. Besides, we run our ablations on the Bigsoundbank split from WavText5k (Deshmukh et al., 2022), which serves as an out-of-distribution evaluation for our models. This is to prevent biasing the evaluation based on the training data. Finally, to further validate our results we run a user preference study. Details about the user study can be found in the Appendix.